Two 6µF capacitance are connected in series with one having a capacitance of 12µF.Find the total equivalent circuit capacitance. What capacitance must be added in series to obtain a capacitance of 1.2µF.
1
Expert's answer
2020-04-28T09:46:18-0400

When three capacitors are connected in series, the equivalent capacitance will be

"C=\frac{1}{1/C_1+1/C_2+1/C_3}=\\
\space\\
=\frac{1}{1/6+1/6+1/12}=2.4\space\mu\text{F}."

Now solve another problem, that is, find what capacitance must be added in series if the total capacitance is 1.2 µF:


"1.2=\frac{1}{1/2.4+1/C_x},\\
\space\\
C_x=\frac{1}{1/1.2-1/2.4}=2.4\space\text{µF}."